he following labor standards have been established for a particular product:

Standard labor hours per unit of output 4.2 hours
Standard labor rate $ 20.00 per hour

The following data pertain to operations concerning the product for the last month:

Actual hours worked 6,800 hours
Actual total labor cost $ 136,680
Actual output 1,600 units

Required:

a. What is the labor rate variance for the month?

b.  What is the labor efficiency variance for the month?

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

a. Labor rate variance
b. Labor efficiency variance
Answer #1

A) Labor rate variance

= (SR-AR) * AH

= [20-(136,680/6800)]*6800

= (20-20.10)*6800

= 680 U

B) Labor efficiency variance

= (SH-AH) *SR

= [(4.2*1600)-6800]*20

= (6720-6800)*20

= 1600 U

Kindly comment if you face any issues